You run a school in Florida. Fixed monthly cost is $5,560.00 for rent and utilities, $5,772.00 is spent in salaries and... HELP WITH THESE TWOYou run a school in Florida. Fixed monthly cost is $5,560.00 for rent and utilities, $5,772.00 is spent in salaries and $1,69
Answer #1

1. Net Profit = Fees from students - Costs

Let the indifference point be x students

For Florida,
Costs = Rent and utilities + salaries + insurance + variable cost per student = 5560 + 5772 + 1685 + 103x =
Fees from Students = 735x
Net Profit = 735x - (5560 + 5772 + 1685 + 103x)

For new neighbourhood,
Costs = Rent and utilities + salaries + insurance + variable cost per student = 10719 + 6605 + 2456 + 157x =
Fees from Students = 1028x
Net Profit = 1028x - (10719 + 6605 + 2456 + 157x)

At indifference point, net profit will be equal for both

=> 735x - (5560 + 5772 + 1685 + 103x) = 1028x - (10719 + 6605 + 2456 + 157x)

=> 735x - 13017 - 103x = 1028x - 19780 - 157x

=> (1028 - 157 - 735 + 103)x = 19780 - 13017

=> x = 6763/239 = 28.29 or 30 students

2.

Let the indifference point be x

Initial pricing -
Selling Price = 176
Variable cost = 21
Rent = 3728
Insurance = 244
Net Profit = Sales - Costs = 176x - (21x + 3728 + 244)

New Pricing
Selling Price = 307
Variable cost = 49
Rent = 6854
Insurance = 401
Net Profit = Sales - Costs = 307x - (49x + 6854 + 401)

At indifference point, the profit would be same

=> 176x - (21x + 3728 + 244) = 307x - (49x + 6854 + 401)

=> 176x - 21x - 3972 = 307x - 49x - 7255

=> (307 - 49 - 176 + 21)x = 7255 - 3972

=> x = 31.87 or 32